OpenStudy (sarahlitts):

PLEASE HELP OH MY GOODNESS! I know the answer, i just need to know how to get it! Will medal and fan! rest of question in comments :) An investment account earns 4% per year compounded annually. If the initial investment was 4,000 dollars, how much is in the account after 3 years?

OpenStudy (sarahlitts):

I know the answer i just need the work! The answer is 4,499

OpenStudy (anonymous):

the answer is correct

OpenStudy (welshfella):

the formula for compound interest is Ammount after n years = P(1 + r/100)^n where r = annual rate per cent, P = principal ( starting amount) and n = number of years

OpenStudy (sarahlitts):

oh my goodness thank you! that helps a ton!

OpenStudy (welshfella):

just plug in your given values

OpenStudy (sarahlitts):

ok ^-^

OpenStudy (welshfella):

yw

OpenStudy (danjs):

just think of it as at the end of each year, you multiply your total current money by 1.04 so you get a string of 1.04's multiplied to your beginning amount , one for each year passed, so it is an exponential current value = (start amt)*(increase)^(years) y = A*b^x

OpenStudy (sarahlitts):

thank you!